PQ = 2x * 1 and QR = 5x - 44 Find PQ
1
Expert's answer
2020-10-08T16:30:03-0400

Given that Q is the midpoint of PR means,

PQ = QR

 

But, PQ = 2x + 1 and QR = 5x - 44

 

Hence we can say that,

2x + 1 = 5x - 44

3x = 45

x = 15

 

Thus, PQ = 2*15 + 1 = 31,

PQ = 31.